OpenStudy (anonymous):

evaluate, ∑(i^2-2i+1) i=1 n=4

OpenStudy (campbell_st):

this is the sum of the terms 1st term i =1 1^2 - 2x1 + 1 2nd term i = 2 2^2 - 2x2 +1 3rd term i = 3 3^2 - 2x3 +1 4th term i = 4 4^2 - 2x4 + 1 you need to evaluate each then add the answers

OpenStudy (anonymous):

hey thanks man ;) this website is great!